OpenStudy (anonymous):

se the substitution method to find all solutions of the system of equations. 2x+ y= 11 x +2y= 4

OpenStudy (anonymous):

use*

OpenStudy (anonymous):

\[y = 11 - 2x\]

OpenStudy (anonymous):

substitute in the second equation \[x + 2(11-2x)=4\]

OpenStudy (anonymous):

solve for x -3x =4-22 x= 6

OpenStudy (anonymous):

go back to first equation \[y= 11- 2(6)\] y= -1